# Scalable Binary CUR Low-Rank Approximation Algorithm

### v1

@@ I-A CUR Low-Rank Approximaation
Let $A\in\mathbb{R}^{m\times n}$ be a given matrix, and consider two sets of indices $I=\{i_{1},\cdots,i_{k}\}$ and $J=\{j_{1},\cdots,j_{k}\}$, corresponding to selected rows and columns, respectively. Define $A_{I,:}$ as the subset of rows in $A$ indexed by $I$, and $A_{:,J}$ as the subset of columns indexed by $J$. Suppose the submatrix $A_{I,J}$, formed by the intersection of rows $I$ and columns $J$, is nonsingular. In this case, the approximation
$A\approx A_{:,J}A_{I,J}^{\dagger}A_{I,:}$
is referred to as a CUR approximiation of $A$. For a visual representation, see Figure 1.
@@
There has been extensive research on CUR Low-Rank Approximations for matrices. For further details, refer to [7, 14, 19, 22, 30, 32, 31] and the references therein. There are different strategies to obtain such an approximations or decomposition in the literature. These include volume optimization [9, 16, 23, 17, 1, 28, 18, 34], rank and spectrum-revealing factorization [20, 3], probabilistic sampling [32, 31, 33, 7, 12, 22], and empirical interpolation [8, 13, 30]. This paper focuses on the first category within the framework of parallel computing.
@@ I-B Maximal Volume Optimization
The CUR approximation with $r$ selected columns and rows is usually formulated as the following optimization problem [9, 16, 23, 17, 1, 28, 18, 34]:
@@
Each iteration in Algorithm 1 identifies a single pivot by searching for the largest absolute value in the residual matrix. Although local maxima searches can be conducted in parallel, a global reduction is necessary to determine the unique pivot element, thus creating an unavoidable sequential bottleneck. Once the pivot is found, the algorithm performs a rank-1 update on the residual, and this updated matrix in turn dictates the location and value of the next pivot. Consequently, there is a strict iteration-by-iteration data dependency, preventing the selection of multiple pivots simultaneously. In this paper, we propose the Scalable Binary CUR Low-Rank Approximation Algorithm. This framework is specifically designed to enable scalable CUR approximations for large-scale matrices.
@@ IV Conclusion
In conclusion, we have presented a Scalable Binary CUR Low-Rank Approximation Algorithm. This approach is designed to overcome the scalability challenges inherent in the CUR decomposition of large-scale matrices in a deterministic way. By leveraging the parallel processing capabilities of modern multi-core architectures, our approach has shown modest improvements in computational efficiency. Future work will focus on refining and optimizing the algorithm to achieve more reliable outcomes in practical data processing scenarios.

### v2

@@ I-A CUR Low-Rank Approximaation
Let $A\in\mathbb{R}^{m\times n}$ be a given matrix, and consider two sets of indices $I\subset[m]$ and $J\subset[n]$, corresponding to selected rows and columns, respectively. Define $A_{I,:}$ as the subset of rows in $A$ indexed by $I$, and $A_{:,J}$ as the subset of columns indexed by $J$. Suppose the submatrix $A_{I,J}$, formed by the intersection of rows $I$ and columns $J$, is nonsingular. In this case, the approximation
$A\approx CUR,$
where $C=A(:,J),U=A^{\dagger}(I,J),R=A(I,:),$ is referred to as a CUR approximiation of $A$. For a visual representation, see Figure 1.
@@
There has been extensive research on CUR Low-Rank Approximations for matrices. For readers interested in further details and additional references, please see the following works [1, 2, 4, 5, 9, 10, 11, 12, 13, 14, 15, 17, 19, 20, 21, 22, 23, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 36, 40, 41, 46, 47, 48, 49, 50, 51, 52, 53] and the references cited therein. There are different strategies to obtain such CUR approximations or decomposition. This paper explore maximal volume optimization strategy within the framework of parallel computing.
@@ I-B Maximal Volume Optimization
The CUR approximation with $r$ selected columns and rows is usually formulated as the following optimization problem [2, 15, 25, 26, 27, 37, 44, 53]
@@
Each iteration in Algorithm 1 identifies a single pivot by searching for the largest absolute value in the residual matrix. Although local maxima searches can be conducted in parallel, a global reduction is necessary to determine the unique pivot element, thus creating an unavoidable sequential bottleneck. Once the pivot is found, the algorithm performs a rank-1 update on the residual, and this updated matrix in turn dictates the location and value of the next pivot. Consequently, there is a strict iteration-by-iteration data dependency, preventing the selection of multiple pivots simultaneously. In this paper, we propose the Scalable Binary CUR Low-Rank Approximation Algorithm. This framework is specifically designed to explore scalable CUR approximations for large-scale matrices by selecting rows and columns in parallel.
@@ IV Conclusion
In conclusion, we have presented a Scalable Binary CUR Low-Rank Approximation Algorithm. This approach is designed to achieve CUR decomposition of large-scale matrices by deterministiclly selecting rows and columns in parallel. By leveraging the parallel processing capabilities of modern multi-core architectures, our approach has shown modest improvements in computational efficiency.